GIF2TIFF(1) GIF2TIFF(1) NAME gif2tiff−createaTIFFfilefromaGIF87formatimagefile SYNOPSIS gif2tiff[options]input.gifoutput.tif DESCRIPTION Gif2tiffconvertsafileintheGIF87formattoTIFF.TheTIFFimageiscreatedasapalette image,withsamplescompressedwiththeLempel-Ziv&Welchalgorithm(Compression=5). Thesecharacteristicscanoverriden,orexplicitlyspecifiedwiththeoptionsdescribedbelow. OPTIONS −cSpecifyacompressionschemetousewhenwritingimagedata:−cnonefornocompression,-cpackbitsforthePackBitscompressionalgorithm,-czipfortheDeflatecom-pressionalgorithm,and−clzwforLempel-Ziv&Welch(thedefault). −rWritedatawithaspecifiednumberofrowsperstrip;bydefaultthenumberofrows/strip isselectedsothateachstripisapproximately8kilobytes. NOTES TheprogramisbasedonPaulHaeberli’sfromgifprogramwhich,inturn,isbasedonMarcelJ.E. Mol’sGIFreader. BUGS Shouldhavemoreoptionstocontroloutputformat. SEEALSO pal2rgb(1),tiffinfo(1),tiffcp(1),tiffmedian(1),libtiff(3) October15,1995 1